OpenStudy (anonymous):

Given that \(\large \color{green}{\int\limits_0^4(x^3 \cdot \sqrt{9 + x^2}).dx = a}\), what is the value of \(\large\color{blue}{[a]}\), where [] is Greatest Integer Function...

OpenStudy (anonymous):

*

OpenStudy (agent0smith):

\[\large \color{green}{\int\limits\limits_0^4(x^3 \cdot \sqrt{9 + x^2}).dx }\] let \[\large x = 3 \tan \theta\] \[\large dx = 3 \sec^2 \theta d \theta\] \[\large \int\limits\limits\limits\limits\limits_0^4(3 \tan \theta)^3 \cdot \sqrt{9 + (3 \tan \theta )^2} (3 \sec^2 \theta d \theta)\]

OpenStudy (anonymous):

Sorry, then??

OpenStudy (anonymous):

Take two..

OpenStudy (agent0smith):

\[\large \int\limits\limits\limits\limits\limits\limits_0^4(3 \tan \theta)^3 \cdot \sqrt{9 + (3 \tan \theta )^2} (3 \sec^2 \theta d \theta) = \] \[\large 27 \times 3 \times \int\limits_{0}^{4} \tan^3 \theta \cdot \sqrt{9(1+ \tan^2 \theta )} \sec^2 \theta d \theta)\] follow to here?

OpenStudy (agent0smith):

\[\large 81 \int\limits\limits_{0}^{4} \tan^3 \theta \cdot \sqrt{9\sec^2 \theta } \sec^2 \theta d \theta) =\] \[\large 243 \int\limits\limits\limits_{0}^{4} \tan^3 \theta \cdot \sec \theta \sec^2 \theta d \theta) =\] \[\large 243 \int\limits\limits\limits\limits_{0}^{4} \tan^3 \theta \cdot \sec^3 \theta d \theta =\]

terenzreignz (terenzreignz):

I never get tired of this... Miss Singh is missing again :3

OpenStudy (agent0smith):

Oh crap i kept the limits 0 to 4, ah well just ignore those till the end I guess.

OpenStudy (agent0smith):

Now use tan^2 = sec^2 - 1\[\large 243 \int\limits\limits\tan^3 \theta \cdot \sec^3 \theta d \theta = 243 \int\limits\limits (\sec^2 \theta -1) \tan \theta \sec^3 \theta d \theta =\] \[\large 243 \int\limits\limits\limits (\sec^2 \theta -1) \tan \theta \sec^3 \theta d \theta = 243 \int\limits \sec^5 \theta \tan \theta - \tan \theta \sec^3 \theta d \theta \] Make yet another substitution (wondering if there was a faster way to do this... by parts?) \[u = \sec \theta \] \[du = \tan \theta \sec \theta d \theta \]

terenzreignz (terenzreignz):

No need o.O \[\large 243\int\limits_{0}^{4}\tan^2\theta\sec^2\theta(\sec\theta\tan\theta)d\theta\]\[\large 243\int\limits_{0}^{4}(\sec^2\theta-1)\sec^2\theta(\sec\theta\tan\theta)d\theta\]

OpenStudy (agent0smith):

And you'll get \[\large \left[ \frac{\sec^5 \theta}{5} - \frac{ \sec^3 \theta} {3}\right]\]

terenzreignz (terenzreignz):

Not yet done. \[\large x=3\tan\theta\]\[\huge \theta=\frac13\tan^{-1}x\]

OpenStudy (agent0smith):

I know :P x=3tanθ and x=0 to x=4

terenzreignz (terenzreignz):

Yeah... but integrating was the easy part. Now, we still have to evaluate it :D

OpenStudy (agent0smith):

If it was an indefinite integral, you can change it back to non-theta-crap with a right triangle  \[\cos \theta = \frac {3} {\sqrt{9+x^2}}\] so \[\large \frac{ 1 }{ \cos \theta } = \sec \theta = \frac {\sqrt{9+x^2}}{3}\] and then put that back into... \[\large \left[ \frac{\sec^5 \theta}{5} - \frac{ \sec^3 \theta} {3}\right] = \] \[\LARGE \left[ \frac{(\sqrt{9+x^2})^5 }{3^5 \times 5} - \frac{ (\sqrt{9+x^2})^3 } {3^3 \times 3}\right]_{0}^{4}\]

OpenStudy (agent0smith):

Oh, throw a 243 in front of that, i think we dropped it. \[\LARGE 243 \left[ \frac{(\sqrt{9+x^2})^5 }{3^5 \times 5} - \frac{ (\sqrt{9+x^2})^3 } {3^3 \times 3}\right]_{0}^{4}\]

OpenStudy (agent0smith):

Which, thank god, agrees with the number wolfram alpha gets! Then you just use the greatest integer function, so you essentially just round down.

OpenStudy (anonymous):

Miss Singh ??? @terenzreignz

terenzreignz (terenzreignz):

Your profile says "Gurpreet Singh" after all. Also, puns ^.^

OpenStudy (anonymous):

What is the idea behind 'Miss' there??

terenzreignz (terenzreignz):

I don't know, you said you were a girl, over chat? Besides, Miss Singh sound so much awesomer and punnier :3

terenzreignz (terenzreignz):

*sounds